OpenStudy (anonymous):

Simplify -3x+9 over -3 A) x-3 B) x+9 C) -x-3 D) -3x-3

OpenStudy (anonymous):

A

OpenStudy (anonymous):

(-3x+9)/(-3)=-3(x-3)/(-3)=x-3

OpenStudy (anonymous):

It's a fraction? Are you guys sure that is right?

OpenStudy (anonymous):

\[-3x/-3=x\] 9/-3=-3 so x-3

OpenStudy (anonymous):

100%

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

I have another one if you guys can help again.

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

1/2(g-16) A) 1/2g-8 B) 1/2g+8 C) 1/2g+16 D) 1/2g-16

OpenStudy (anonymous):

1/2=0.5 0.5(g-16)=0.5g-8=(1/2)g-8

OpenStudy (anonymous):

The fisr one

OpenStudy (kymber):

You need to distribute the 1/2